a man walk 3 mph. he covers 13.5 miles how long does it take?
also a cyclist travles of 8 m/s. how far will she travel ?AND
A bus travels between two towns 16 kilometers apart. it take 30 minutes.what is the average speed.
From luca:P

Answer by DrBeeee(684)

It is standard practice to use the following variables in distance problems
(1) d = distance
(2) r = rate
(3) t = time
Us math wizards however, must be careful that the proper units of these three variables are consistant. Don't worry, I'll show you on each example.
There are three forms of the formula:
(4) d = r*t
(5) r = d/t and
(6) t = d/r
Now let's do your problems.
In (a) you want the time so use (6) to get
(7) t = d/r
(8) t = 13.5/3 or
(9) t = 4.5 hr
I'll finish later, gotta go to soccer game